Let X ? Poisson (?).

a. Show that the Poisson probabilitiesP= P(X = ) for x = 0, 1, 2, . . . can be computed recursively by
p_0 = e -\lambda andp_k=\frac{\lambda}{k}p_{k-1} for k = 1, 2, . . ..

b. Use the scheme to find P(X ? 4) for ? = 4.5.
Answer #1

a)

We know that p_{x}=P(X=x)=\frac{\lambda^{x}e^{-\lambda}}{x!}

p_{0}=e^{-\lambda}

Now, p_{1}=\frac{\lambda}{1}p_{0}=\lambda e^{-\lambda}

and p_{2}=\frac{\lambda}{2}p_{1}=\frac{\lambda}{2}(\lambda e^{-\lambda})=\frac{\lambda^{2}e^{-\lambda}}{2!} and so on...

So, we conclude that Poisson probabilities can be computed recursively.

b)

Now, p_{0}=e^{-\lambda}=e^{-4.5}

p_{1}=\frac{\lambda}{1}p_{0}=4.5 e^{-4.5}

p_{2}=\frac{\lambda}{2}p_{1}=\frac{\lambda}{2}(\lambda e^{-\lambda})=\frac{4.5^{2}e^{-4.5}}{2!}

p_{3}=\frac{\lambda^{3}e^{-\lambda}}{3!}=\frac{4.5^{3}e^{-4.5}}{3!}

p_{4}=\frac{\lambda^{4}e^{-\lambda}}{4!}=\frac{4.5^{4}e^{-4.5}}{4!}

P(X\leq 4)=p_{0}+p_{1}+p_{2}+p_{3}+p_{4}

=e^{-4.5}+\frac{4.5^{1}e^{-4.5}}{1!}+\frac{4.5^{2}e^{-4.5}}{2!}+\frac{4.5^{3}e^{-4.5}}{3!}+\frac{4.5^{4}e^{-4.5}}{4!}=0.5321